OpenStudy (anonymous):

Put the following radical expression into simplified form. Assume the variable represents a positive number.

OpenStudy (anonymous):

\[\sqrt{\frac{ 98a^5 }{ 3 }}\]

hartnn (hartnn):

factor 98

OpenStudy (anonymous):

49 & 2

OpenStudy (anonymous):

49=7

hartnn (hartnn):

yes what about \(\sqrt {a^5}\) caqn you now simplify that ? after practicing so many similar problems ?

OpenStudy (anonymous):

a^4*a=a^2=a \[a \sqrt{a}\]

hartnn (hartnn):

sure ?

OpenStudy (anonymous):

yea, cuz one a will be left inside the radical, right?

Miracrown (miracrown):

lets try simplifying the 98 first we are dealing with a sqrt so that means we can pull out multiples of two of a kind so lets break down 98 into its multiples

hartnn (hartnn):

\(\Large \sqrt {a^5} = \sqrt {a^4}\sqrt a = a^2 \sqrt a\)

hartnn (hartnn):

got that^^ ?

OpenStudy (anonymous):

i thought the a^2 would still be in radical, that's why

Miracrown (miracrown):

scratch what I said >.< Hartnn is already helpin

OpenStudy (anonymous):

yup!

hartnn (hartnn):

cool, so you have your numerator now what about denominator ? how will you rationalize it ? (once again, you've done it many times :) )

OpenStudy (anonymous):

multiply it & it becomes 3 w/o the radical

hartnn (hartnn):

yes, so final answer will be ?

OpenStudy (anonymous):

Always tempted

OpenStudy (anonymous):

\[\frac{ 7a^2\sqrt{2} }{ 3 }\]

OpenStudy (anonymous):

2a

hartnn (hartnn):

2a yes, correct! :)

hartnn (hartnn):

wait, there's sqrt 3 too!

hartnn (hartnn):

so sqrt (6a)

hartnn (hartnn):

right ?

OpenStudy (anonymous):

I don't know, 6a where?

hartnn (hartnn):